OpenStudy (anonymous):

50% off of $75.00 = $37.50 What is 75% off of $75.00?

OpenStudy (lgbasallote):

do you know what 75% is in decimal?

OpenStudy (anonymous):

.75

OpenStudy (lgbasallote):

multiply that to $75.00

OpenStudy (anonymous):

56.25 - I understand how that is the answer but I feel like it should be less than 37.50 (which is 50% off of $75)

OpenStudy (lgbasallote):

which is bigger..75 or 50? /:)

OpenStudy (anonymous):

I'm going to give you the original question even though its long. a. Suppose you purchase an item from the Trim-A-Tree shop for $75.00 (original price). What is the sale price? My answer: The sale price of an item from the Trim-A-Tree shop is $37.50. b. Senior citizens (age 62 and older) were eligible for a senior discount. Assume someone over age 62 purchased the same item on sale (originally $75.00) and received an additional 15% off the sale price. What price do senior citizens pay for the same item from the Trim-A Tree Shop?

OpenStudy (anonymous):

shouldn't 75% off of an item have a smaller price rather than 50% off?

OpenStudy (lgbasallote):

ugh no wonder....discounts *facepalm* yes it should be smaller...and it shouldn't be multiplied by 75% it should be multiplied by 25% a shortcut would be \(100% - \text{given discount rate}) \times \text{original price}\) since given discount rate is 75% and original price is 75$ we plug them in \(100% - 75% \times 75 = 25% \times 75\) got it?

OpenStudy (lgbasallote):

uhmm did the latex become visible there ^? or is it only showing 100?

OpenStudy (anonymous):

lol its just showing 100

OpenStudy (anonymous):

so it should be .25*75.00 = 18.75 right?

OpenStudy (lgbasallote):

yupp anyway the formula i wrote was (100% - given discount rate) x original price

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

56.25 37.50/2=18.75 18.75+37.50=75%=56.25

OpenStudy (anonymous):

^ That would be right if it was not a discount but it is, thank you anyway!